Solve the following linear systems by the
process of elimination : 3
x + 2y + 5z = 9
x — y + $z = 2
3x — 6y z = 26
1
Expert's answer
2019-11-26T14:26:01-0500

3x + 2y + 5z = 9

x - y + z = 2

3x -6y - z = 26



3x +2y +5z = 9

x =2 + y - z

3x -6y -z = 26



3(2 +y -z) + 2y +5z = 9

x =2 + y - z

3(2+y-z) -6y -z=26



5y + 2z = 3 |•2

x =2 + y - z

-3y - 4z = 20



10y + 4z = 6

x =2 + y - z

4z = -3y -20



10y -3y -20 = 6

x =2 + y - z

4z = -3y -20



7y = 26

x =2 + y - z

4z = -3y -20



y=26/7

x =2 + y - z

4z = -3y -20



y=26/7

x=2+26/7 - z

4z=-20-78/7



y=26/7

x=40/7 - z

z=-109/14



y=26/7

x=189/14

z=-109/14